Alexa owns a Coffee Shop. The following information is from her accountant:

Expenses:

Advertising $55.00

Coffee, cups, sweetener, flour, etc. $9.27 per cup of coffee and a muffin

Monthly rental $126.00

What will her break-even point be if she sets the price for coffee and a muffin at $13.27?

** Round to 0 Decimal Places! **
Answer #1

Variable expense per unit = $9.27

Selling price per unit = $13.27

Contribution per unit = Selling price - Variable cost

= 13.27 - 9.27

= $4.00

Fixed cost = advertising + monthly rental

= 55 + 126

= 181

Break even point (in units) = Fixed costs / Contribution per unit

= 181 / 4

= 45.25

Or 45